### v1

# Functional Stochastic Gradient MCMC for Bayesian Neural Networks

Abstract
Classical variational inference for Bayesian neural networks (BNNs) in parameter space usually suffers from unresolved prior issues such as knowledge encoding intractability and pathological behaviors in deep networks, which could lead to an improper posterior inference. Hence, functional variational inference has been proposed recently to resolve these issues via stochastic process priors. Beyond variational inference, stochastic gradient Markov Chain Monte Carlo (SGMCMC) is another scalable and effective inference method for BNNs to asymptotically generate samples from true posterior by simulating a continuous dynamic. However, the existing SGMCMC methods only work in parametric space, which has the same issues of parameter-space variational inference, and extending the parameter-space dynamics to function-space dynamics is not a trivial undertaking. In this paper, we introduce a new functional SGMCMC scheme via newly designed diffusion dynamics, which can incorporate more informative functional priors. Moreover, we prove that the stationary distribution of these functional dynamics is the target posterior distribution over functions. We demonstrate better performance in both accuracy and uncertainty quantification of our functional SGMCMC on several tasks compared with naive SGMCMC and functional variational inference methods.
Machine Learning, ICML
1 Introduction
Compared to regular deep neural networks that rely on a single pattern of model parameters only, Bayesian neural networks (BNNs) have demonstrated the ability to provide a principled predictive uncertainty estimation and enhanced generalization performance through a model integration under posterior distribution over network weights (Neal, 1995; Gal et al., 2016; Wilson & Izmailov, 2020). With the continued development and success of Bayesian inference methods in the deep learning community, BNNs have been applied in a wide variety of domains and tasks, including prediction in astrophysics (Cranmer et al., 2021), medical diagnosis and diabetic detection (Filos et al., 2019; Band et al., 2022), and other safety-critical environments (Rudner et al., 2022).
Prior choosing is apparently crucial for BNNs. The existing common practice is i.i.d. Gaussian priors for all network weights, which is also named parameter-space prior. However, this parameter-space prior has been shown to have some problematic issues. For example, as network depth increases, the function samples of Gaussian prior over weights tend to be horizontal (Duvenaud et al., 2014; Matthews et al., 2018), and the prior distribution of unit outputs becomes more heavy-tailed with ReLU activation (Vladimirova et al., 2019; Tran et al., 2022), which will lead to improper posterior inference (Papamarkou et al., 2022). Moreover, the effects of this parameter space prior to the posterior inference remain an open question due to the complex network architecture (Fortuin et al., 2021; Wild et al., 2022), so it is difficult to transform the valid prior knowledge into corresponding network weights’ prior. Given these unresolved prior problems for parameters, Bayesian inference directly in function space has gained increasing interest in recent years since one can incorporate more informative stochastic processes as the prior. For example, Gaussian processes (GPs) which can easily encode the prior knowledge about the potential functions via kernel functions choosing have been widely used as the prior in function-space variational inference for BNNs (Sun et al., 2019; Ma & Hernández-Lobato, 2021; Rudner et al., 2022; Pielok et al., 2022) which has been shown to yield better prediction and improved uncertainty estimation performance than variational inference in parameter space.
For the posterior inference given prior, Markov Chain Monte Carlo (MCMC) is another valid method for BNNs apart from the commonly used variational inference (Blundell et al., 2015; Sun et al., 2019). Instead of optimizing a variational posterior by minimizing the distance (divergence) with the true posterior (Graves, 2011; Blundell et al., 2015; Gal & Ghahramani, 2016), MCMC approaches are regarded as the gold standard for Bayesian inference since they generate discrete samples from a continuous-time Markov chain whose stationary distribution is exactly the true posterior (Neal et al., 2011; Welling & Teh, 2011; Li et al., 2016). So, they do not need to make strong assumptions about the distributional form of the posterior (Alexos et al., 2022; Cobb & Jalaian, 2021). Even better, the computationally expensive MCMC methods became feasible for modern BNNs with large data following the introduction of their scalable stochastic gradient variants (SGMCMC) (Welling & Teh, 2011; Patterson & Teh, 2013; Chen et al., 2014; Zhang et al., 2020; Nemeth & Fearnhead, 2021). However, all existing MCMC methods for BNNs are still in parameter space, and extending them to function space is not a minor process because we need to design novel dynamics in function space and ensure the new dynamics would have the target posterior as the stationary distribution.
In this work, we propose a functional SGMCMC scheme based on the fundamental Itô lemma for stochastic calculus (Itô, 1951; Brzeźniak et al., 2008; Øksendal, 2013). The dynamics of BNNs in function space are designed by proposing a new potential energy function where we integrate the more informative functional prior, such as GPs, into the diffusion process. Furthermore, we prove that the stationary distribution of the new functional dynamics is the desired posterior distribution over functions. Our main contributions are as follows:
  • We propose a new functional SGMCMC approach for BNNs that can utilize the more informative functional prior to the posterior inference.
  • We prove that under our designed functional SGMCMC dynamics, the stationary distribution would be the target posterior distribution over functions of BNNs.
  • Our method shows improved predictive performance and uncertainty estimation in several benchmark tasks compared to parameter-space SGMCMC and parameter- and function-space variational inference methods.
2 Preliminaries
2.1 Bayesian Neural Networks
[…]
2.2 Stochastic Gradient MCMC for BNNs
Traditional MCMC methods, such as the random-walk Metropolis-Hastings (MH) algorithm (Metropolis et al., 1953), are easy to apply but suffer from a slow mixing rate and high correlation between samples. In contrast, the dynamics-based sampling methods utilizing gradients of log posterior density can provide a more efficient exploration of the posterior (Girolami & Calderhead, 2011). Also, their stochastic gradient variants (Welling & Teh, 2011; Chen et al., 2014; Baker et al., 2019), such as the stochastic gradient Langevin dynamics (SGLD) and the stochastic gradient Hamiltonian Monte Carlo (SGHMC), can scale to large datasets for practical applications.
SGLD. The objective of MCMC is to obtain random samples of the posterior distribution of model parameters (with the unnormalized form) $p(\mathbf{w}|\mathcal{D})\propto\exp(-U(\mathbf{w}))$, where $U(\mathbf{w})=-\log p(\mathbf{Y_{\mathcal{D}}}|\mathbf{X_{\mathcal{D}}},\mathbf{w})-\log p(\mathbf{w})$ is the potential energy function. The idea of dynamics-based sampling for $p(\mathbf{w}|\mathcal{D})$ is to simulate a continuous-time stochastic process whose stationary distribution is the target posterior. Combining Langevin dynamics (Neal et al., 2011) with stochastic Robbins-Monro optimization algorithm (Robbins & Monro, 1951), Welling & Teh (2011) proposed SGLD for scalable MCMC algorithms. The approximated discretization transition rule for posterior samples is
$\mathbf{w}_{t+1}=\mathbf{w}_{t}-\epsilon_{t}\nabla\tilde{U}(\mathbf{w}_{t})+\sqrt{2\epsilon_{t}}\eta_{t},$ (1)
where $\epsilon_{t}$ is the step-size, $\eta_{t}$ is a standard Gaussian, and $\nabla\tilde{U}(\mathbf{w})$ uses mini-batch stochastic gradient for the likelihood term as $\nabla\tilde{U}(\mathbf{w})=-\frac{N}{n}\sum_{i=1}^{n}\nabla\log p(y_{i}|x_{i},\mathbf{w})-\nabla\log p(\mathbf{w})$ and $n\ll N$. With the decreasing step-size $\epsilon_{t}$, the stochastic gradient noise tends to zero faster than the injected Gaussian noise $\eta_{t}$, and the discretization error is eliminated, then the MH step can be ignored.
SGHMC. To enable a more efficient exploration in the state space, Hamiltonian dynamics (Duane et al., 1987; Neal, 2012) introduces a momentum variable $\mathbf{z}\in\mathbb{R}^{d}$, and the new joint distribution is $p(\mathbf{w}|\mathcal{D})p(\mathbf{z})\propto\exp(-H(\mathbf{w},\mathbf{z}))$ with corresponding energy function $H(\mathbf{w},\mathbf{z})=-\log p(\mathbf{Y_{\mathcal{D}}}|\mathbf{X_{\mathcal{D}}},\mathbf{w})-\log p(\mathbf{w})-\log p(\mathbf{z})$, where $p(\mathbf{z})$ is commonly assumed as a zero-mean Gaussian $\mathcal{N}(\mathbf{z}|0,M)$. The sample rule in the full-batch Hamiltonian Monte Carlo is defined as following with an MH step (Neal, 2012; Cobb & Jalaian, 2021):
$\mathbf{w}_{t+1}$ $=\mathbf{w}_{t}+\epsilon_{t}M^{-1}\mathbf{z}_{t},$ (2)
$\mathbf{z}_{t+1}$ $=\mathbf{z}_{t}-\epsilon_{t}\nabla U\left(\mathbf{w}_{t}\right).$
Like SGLD, SGHMC uses stochastic gradient $\nabla\tilde{U}(\mathbf{w})$ for the sampling of momentum $\mathbf{z}$, also with an additional friction term $\epsilon_{t}M^{-1}\mathbf{z}_{t}$ and a Gaussian injected noise to guarantee the (marginal) stationary distribution of the Hamiltonian dynamics still be the target posterior distribution (Chen et al., 2014). The specific update rule for samples of $\mathbf{z}$ is as: $\mathbf{z}_{t+1}=\mathbf{z}_{t}-\epsilon_{t}\nabla\tilde{U}(\mathbf{w}_{t})-\epsilon_{t}CM^{-1}\mathbf{z}_{t}+\sqrt{2C\epsilon_{t}}\eta_{t}$, where $C>0$ is a scalar coefficient.
3 Functional Stochastic Gradient MCMC
Generally, dynamics-based MCMC approaches are derived from the Itô diffusion (Øksendal, 2003), which is similar to the Brownian motion of the particle subjected to a potential in fluid physics. In this section, to solve the parameter-space prior issues in existing SGMCMC methods, we propose a new functional SGMCMC scheme via a newly designed diffusion dynamics, which is capable of incorporating more informative functional priors. In Section 3.1, we first review the Itô diffusion for dynamics-based MCMC. Then we design a new functional Langevin dynamics directly for the function mapping defined by the BNNs. We prove that the stationary distribution of such functional Langevin dynamics is the target posterior distribution over functions. Furthermore, in Section 3.2, we design a functional Hamiltonian dynamics and the corresponding stochastic gradient sampling scheme for BNNs. We also prove it can guarantee the desired invariant distribution.
3.1 Functional SGLD for BNNs
An Itô diffusion in n-dimensional Euclidean space is defined as a stochastic differential equation (SDE) to describe the evolution of a continuous-time stochastic process $X:[0,\infty)\times\Omega\rightarrow\mathbb{R}^{n}$ on a probability space $(\Omega,\Sigma,P)$. The general form of an Itô diffusion driven by the standard Wiener process (Brownian motion) is as follows:
$\mathrm{d}X_{t}=\mu(X_{t})\mathrm{d}t+\sigma(X_{t})\mathrm{d}B_{t},$ (3)
where $X_{t}$ represents the state of the stochastic process at time $t$, $B\in\mathbb{R}^{n}$ is an n-dimensional Brownian motion, $\mu(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is a vector field denoting the deterministic drift term and matrix field $\sigma(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times n}$ is known as the diffusion coefficient for $X_{t}$. Both $\mu(\cdot)$ and $\sigma(\cdot)$ are assumed to satisfy the usual Lipschitz continuity condition (Ghosh, 2010). And $\mathrm{d}B_{t}$ is the increment of a Wiener process distributed as $\mathrm{d}B_{t}\sim\mathcal{N}(0,\mathrm{d}t\cdot\mathbf{I}_{n})$. Let $p(x,t)$ be the probability density of random variable $X_{t}$, and the corresponding evolution of this probability density of the diffusion process can be described by the Fokker–Planck equation (Risken, 1996), which is the key to derive the dynamics-based MCMC (see Appendix B).
Naturally, the idea of existing dynamics-based SGMCMC methods for BNNs is to simulate a continuous diffusion of network weight whose stationary distribution is the target posterior $p(\mathbf{w}|\mathcal{D})$ with appropriately designed drift term and diffusion coefficient. To incorporate informative functional prior into the posterior inference process, we now directly consider the dynamics of function mapping $f(\cdot;\mathbf{w})$ defined by BNNs. According to Itô Lemma (Itô, 1951), $f(\cdot;\mathbf{w})$ is itself an Itô diffusion (Brzeźniak et al., 2008) (see Appendix A). That is, for exiting SGMCMC methods, the stochastic process $X$ simulated in Itô diffusion is usually for network weight or its augmentation, which we now lift to the stochastic function mapping of BNNs directly as
$\mathrm{d}f_{t}=\mu(f_{t})\mathrm{d}t+\sigma(f_{t})\mathrm{d}B_{t}.$ (4)
In a probability space $(\Omega,\mathcal{F},P)$ (separable and complete Polish space), suppose $p(f)$ is the prior distribution over functions, such as the classical Gaussian process (GP) prior that can easily encode prior knowledge about function properties. $p(\mathbf{Y_{\mathcal{D}}}|f(\mathbf{X_{\mathcal{D}}}))$ is the likelihood function with observation $\mathcal{D}=\left\{\mathbf{X_{\mathcal{D}}},\mathbf{Y_{\mathcal{D}}}\right\}$ and $p(f|\mathcal{D})$ is the posterior over functions. For brevity, we abbreviate them as $p(f)$, $p(\mathcal{D}|f)$ and $p(f|\mathcal{D})$, respectively. Let $U(f)=-\log p(\mathcal{D}|f)-\log\frac{\mathrm{d}p}{\mathrm{d}r_{\mathcal{H,F}}}(f))$ be the potential energy functional, where $r_{\mathcal{H,F}}$ is the corresponding Hausdorff / Riemannian measure, and $\mu(f)=-\nabla_{f}U(f),\sigma(f)=\sqrt{2}\cdot\mathbf{I}_{\infty}$ for the function mapping $f$ of BNNs, we have the following functional Langevin dynamics:
$\mathrm{d}f_{t}=-\nabla_{f}U(f_{t})\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t},$ (5)
and it is easy to verify that the stationary distribution for $p(f,t)$ is exactly the target posterior distribution over functions as $\pi(f)=\exp(-U(f))=p(f|\mathcal{D})$.
However, even for the finite input case, it is intractable to discretize and sample the above diffusion of function mapping in practice directly. Considering that a function mapping $f(\cdot;\mathbf{w})$ of BNNs is completely determined by the network weights $\mathbf{w}$ (assuming the network architecture is fixed), we solve this issue by transforming the functional Langevin dynamics to a weight-space dynamics of which stationary distribution is still $p(f|\mathcal{D})$. We firstly design the following specific Riemannian functional Langevin dynamics for $f(\cdot;\mathbf{w})$ as follows:
$\mathrm{d}f_{t}(\cdot;\mathbf{w})=$ $\mu(f_{t}(\cdot;\mathbf{w}))\mathrm{d}t+\sigma(f_{t}(\cdot;\mathbf{w}))\mathrm{d}B_{t}$ (6)
$=$ $\left[-(\nabla_{\mathbf{w}}f)^{T}(\nabla_{\mathbf{w}}f)\nabla_{f}U(f_{t})+H_{\mathbf{w}}f\right]\mathrm{d}t$
$+\sqrt{2}(\nabla_{\mathbf{w}}f)^{T}\mathrm{d}B_{t},$
where $\nabla_{\mathbf{w}}f$ is the gradient of $f(\cdot;\mathbf{w})$ w.r.t. weight $\mathbf{w}$, $H_{\mathbf{w}}f$ is the Hessian matrix of $f(\cdot;\mathbf{w})$ w.r.t. $\mathbf{w}$. The stationary distribution of this dynamics is guaranteed to be $p(f|\mathcal{D})$ by the following Proposition.
Proposition 3.1.
The stationary distribution of the functional Langevin dynamics defined in Equation 6 is the target posterior distribution over functions $p(f|\mathcal{D})$.
Proof.
The proof is given in Appendix B. ∎
Secondly, we achieve the function-space to weight-space transformation based on the relationship between the dynamics of a random variable with the dynamics of its function. According to the Itô Lemma (Itô, 1951), given the general diffusion framework for weights as $\mathrm{d}\mathbf{w}_{t}=\mu(\mathbf{w}_{t})\mathrm{d}t+\sigma(\mathbf{w}_{t})\mathrm{d}B_{t}$, the differential of the function mapping $f(\cdot;\mathbf{w})$ can be denoted as
$\mathrm{d}f_{t}(\cdot;\mathbf{w})$ (7)
$=(\nabla_{\mathbf{w}}f)^{T}\mathrm{d}\mathbf{w}_{t}+\frac{1}{2}(\mathrm{d}\mathbf{w}_{t})^{T}(H_{\mathbf{w}}f)\mathrm{d}\mathbf{w}_{t}$
$=\left((\nabla_{\mathbf{w}}f)^{T}\mu(\mathbf{w}_{t})+\frac{1}{2}Tr[\sigma(\mathbf{w}_{t})^{T}(H_{\mathbf{w}}f)\sigma(\mathbf{w}_{t})]\right)\mathrm{d}t$
$+(\nabla_{\mathbf{w}}f)^{T}\sigma(\mathbf{w}_{t})\mathrm{d}B_{t}.$
Given our defined functional dynamics in Equation 6 and the above relationship, the corresponding diffusion process of network weights in parameter space under our functional Langevin dynamics can be derived as
$\mathrm{d}\mathbf{w}_{t}$ $=\mu(\mathbf{w}_{t})\mathrm{d}t+\sigma(\mathbf{w}_{t})\mathrm{d}B_{t}$ (8)
$=\nabla_{\mathbf{w}}f[\nabla_{f}(\log p(\mathcal{D}|f_{t})+\log\frac{\mathrm{d}p_{0}}{\mathrm{d}r_{\mathcal{H,F}}}(f))]\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t},$
where the corresponding drift term $\mu(\mathbf{w}_{t})=-\nabla_{\mathbf{w}}U(f_{t})$, and the diffusion coefficient $\sigma(\mathbf{w}_{t})=\sqrt{2}\cdot\mathbf{I}_{d}$. Then, the function samples of the target posterior $p(f|\mathcal{D})$ as the stationary distribution of the functional Langevin dynamics can be obtained by discretizing and sampling the diffusion process of weights in Equation 8.
In practice, the stochastic gradient variant drift term of SDE in Equation 8 is denoted as $-\nabla_{\mathbf{w}}\tilde{U}(f_{t})$, which uses the minibatch version for the likelihood term. For the gradient of the log prior distribution over functions, we firstly assume an approximated GP prior for BNNs, since in the limit of infinite width, the prior measure of BNNs converges to a GP with NNGP kernel (Neal, 1995; Lee et al., 2017; Matthews et al., 2018; Lee et al., 2019). Due to the infinite-dimensional property of functional distribution, we solve the gradient of functional prior on finite measurement points $\mathbf{X}_{\mathcal{M}}\stackrel{{\scriptstyle\text{ dif }}}{{=}}\left[\mathbf{x}_{1},\ldots,\mathbf{x}_{M}\right]^{\mathrm{T}}$ as $\nabla_{f^{\mathbf{X}_{\mathcal{M}}}}\log p(f^{\mathbf{X}_{\mathcal{M}}})$, where $f^{\mathbf{X}_{\mathcal{M}}}$ are corresponding function values evaluated at $\mathbf{X}_{\mathcal{M}}$. Then the prior $p(f^{\mathbf{X}_{\mathcal{M}}})$ is a multivariate Gaussian that could be solved analytically for the gradient. The specific discretization update rule for weight samples under functional stochastic gradient Langevin dynamics (fSGLD) is
$\mathbf{w}_{t+1}=$ $\mathbf{w}_{t}-\epsilon_{t}\nabla_{\mathbf{w}}f\Big{[}-\frac{N}{n}\sum_{i=1}^{n}\nabla_{f}\log p(y_{i}|f_{t}(x_{i};\mathbf{w}_{t}))-$ (9)
$\hskip 62.59596pt\nabla_{f^{\mathbf{X}_{\mathcal{M}}}}\log p(f_{t}^{\mathbf{X}_{\mathcal{M}}})\Big{]}+\sqrt{2\epsilon_{t}}\eta_{t}.$
Similar to the naive SGLD, with the discretization step-size $\epsilon_{t}$ decays to zero, the stationary distribution for functional $p(f,t)$ of fSGLD continues to be the target posterior $p(f|\mathcal{D})$ over functions.
Note that although the function samples of posterior $p(f|\mathcal{D})$ are drawn by the discretization process for corresponding diffusion of weights in Equation 9, our fSGLD is different from the existing SGLD in parameter space, which means the stationary distributions from two dynamics would be different as well. Compared to the naive parameter-space SGLD in Equation 1 controlled only by the limited parameter information, the discretization update rule for weight samples in Equation 9 under our fSGLD scheme contains information about the function mapping, which is important for accurate posterior inference. The pseudocode for fSGLD is shown in Algorithm 1 in Appendix C.
3.2 Functional SGHMC for BNNs
For the naive Hamiltonian dynamics, $X=(\mathbf{w},\mathbf{z})$ is the augmentation of model parameters. We now let $X=(f,g)$ and define the Hamiltonian dynamics in function space as $H(f,g)=U(f)-\log p(g)$, where $p(g)$ is the functional auxiliary distribution induced by the distribution of auxiliary variable $\mathbf{z}$. We design the functional Hamiltonian dynamics for $f(\cdot;\mathbf{w})$ and $g(\cdot;\mathbf{z})$ as
$\mathrm{d}f_{t}(\cdot;\mathbf{w})=-(\nabla_{\mathbf{w}}f)^{T}\nabla_{\mathbf{z}}g\nabla_{g}\log p(g_{t})\mathrm{d}t,$ (10)
$\mathrm{d}g_{t}(\cdot;\mathbf{z})=-(\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{w}}f\nabla_{f}U(f_{t})\mathrm{d}t.$
Table 1: Drift term and diffusion coefficient for naive SGMCMC and functional SGMCMC algorithms.
Dynamics $\mu(X)$ $\sigma(X)$
SGLD $-\nabla\tilde{U}(\mathbf{w})$ $\sqrt{2}\cdot\mathbf{I}_{d}$
fSGLD $-(\nabla_{\mathbf{w}}f)^{T}(\nabla_{\mathbf{w}}f)\nabla_{f}\tilde{U}(f)+H_{\mathbf{w}}f$ $\sqrt{2}(\nabla_{\mathbf{w}}f)^{T}$
SGHMC $-\left(\begin{array}[]{cc}\mathbf{0}&-\mathbf{I}_{d}\\ \mathbf{I}_{d}&C\end{array}\right)\left(\begin{array}[]{cc}\nabla\tilde{U}(\mathbf{w})\\ M^{-1}\mathbf{z}\end{array}\right)$ $\left(\begin{array}[]{cc}\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\sqrt{2C}\end{array}\right)$
fSGHMC $-\left(\begin{array}[]{cc}\mathbf{0}&-(\nabla_{\mathbf{w}}f)^{T}\nabla_{\mathbf{z}}g\\ (\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{w}}f&C(\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{z}}g\end{array}\right)\left(\begin{array}[]{c}\nabla_{f}\tilde{U}(f)\\ -\nabla_{g}\log p(g)\end{array}\right)$ $\left(\begin{array}[]{cc}\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\sqrt{2C}(\nabla_{\mathbf{z}}g)^{T}\end{array}\right)$
Proposition 3.2.
The (marginal) stationary distribution of the functional Hamiltonian dynamics defined in Equation 10 is the target posterior distribution over functions $p(f|\mathcal{D})$.
Proof.
The proof is given in Appendix B. ∎
Similar to the fSGLD, we also transform the above function-space Hamiltonian dynamics to its corresponding weight-space dynamics of $\mathbf{w}$ and $\mathbf{z}$ as follows:
$\mathrm{d}\mathbf{w}_{t}$ $=\frac{\partial H}{\partial\mathbf{z}}=-\nabla_{\mathbf{z}}g\cdot\nabla_{g}\log p(g_{t})\mathrm{d}t,$ (11)
$\mathrm{d}\mathbf{z}_{t}$ $=-\frac{\partial H}{\partial\mathbf{w}}=-\nabla_{\mathbf{w}}f\cdot\nabla_{f}U(f_{t})\mathrm{d}t,$
where the random diffusion term is 0 for the full-batch version.
For the stochastic gradient version, the discretization update rule for samples of $\mathbf{w}$ and $\mathbf{z}$ under the functional stochastic gradient Hamiltonian dynamics (fSGHMC) is as follows:
$\mathbf{w}_{t+1}=$ $~{}\mathbf{w}_{t}-\epsilon_{t}\nabla_{\mathbf{z}}g\cdot\nabla_{g}\log p(g_{t})$ (12)
$\mathbf{z}_{t+1}=$ $~{}\mathbf{z}_{t}-\epsilon_{t}\nabla_{\mathbf{w}}f\cdot\nabla_{f}\tilde{U}(f_{t})$
$\hskip 14.22636pt+\epsilon_{t}C\nabla_{\mathbf{z}}g\cdot\nabla_{g}\log p(g_{t})+\sqrt{2C\epsilon_{t}}\eta_{t}.$
The stationary distribution of such functional stochastic gradient Hamiltonian is still the target functional posterior $p(f|\mathcal{D})$ (see Appendix B). Like the discussion in the last section, the diffusion process for $\mathbf{w}$ and $\mathbf{z}$ under fSGHMC scheme in Equation 12 containing abundant functional information is obviously different from the naive parameter-space SGHMC in Equation 2. The pseudocode for fSGHMC is shown in Algorithm 2 in Appendix C.
For our functional SGMCMC, the predictive distribution is obtained from the following integration process:
$p(y^{*}|x^{*},\mathcal{D})$ $=\int p(y^{*}|f(x^{*}))p(f|\mathcal{D})\mathrm{d}f$ (13)
$\approx\frac{1}{S}\sum_{j=1}^{S}p(y^{*}|f(x^{*};\mathbf{w}^{(j)})),$
where function draws of posterior over functions are generated from corresponding samples of weight $\mathbf{w}^{(j)}(j=1,2,...,S)$ using fSGLD or fSGHMC.
The comparisons of the definitions for the drift term and the diffusion coefficient of popular exiting SGMCMC algorithms and our functional fSGLD and fSGHMC are summarized in Table 1.
4 Related Work
Functional Variational Inference for BNNs. With unresolved concerns about parameter-space weight prior to posterior approximation in BNNs, Bayesian inference in function space has attracted increasing attention. In order to specify meaningful functional prior like Gaussian processes, Sun et al. (2019) proposed a kind of functional variational Bayesian neural networks by maximizing the functional evidence lower bound (ELBO) to optimize the Kullback–Leibler (KL) divergence between the posterior processes and variational posterior over functions explicitly. They solved the KL divergence between infinite-dimensional stochastic processes on a finite marginal measurement set and used the Spectral Stein gradient estimator for implicit functional prior and variational posterior over functions of BNNs (Shi et al., 2018). However, Burt et al. (2020) pointed out some limitations of employing KL divergence in infinite-dimensional function space. The variational posterior should be absolutely continuous with respect to the prior measure to guarantee the existence of the Radon-Nikodym derivatives involved in the definition of functional KL divergence (Matthews et al., 2016; Wild et al., 2022). Otherwise, for example, the KL divergence between distributions over functions generated from two BNNs with different network architectures is ill-defined and could be infinite (Ma & Hernández-Lobato, 2021). To avoid this issue, Rudner et al. (2022) proposed to approximate the distributions over functions of a BNN prior and variational posterior that share the same network structures as two Gaussian distributions via the linearization of their mean parameters and derived a tractable and well-defined functional variational objective for BNNs. Considering the potential weaknesses of KL divergence for stochastic processes, Tran et al. (2022) proposed to match a BNN prior to a GP prior by minimizing the 1-Wasserstein distance (Kantorovich, 1960) to obtain a more interpretable functional prior for BNNs.
Markov Chain Monte Carlo for BNNs. Papamarkou et al. (2022) summarized the challenges in applying MCMC approaches to posterior inference for BNNs in four aspects: computational cost, model structure, weight symmetries, and prior specification. MCMC methods have not been broadly adopted in posterior inference for large modern BNNs until the introduction of scalable Monte Carlo algorithms, such as the stochastic gradient MCMC (Welling & Teh, 2011; Chen et al., 2014; Ahn et al., 2014; Garriga-Alonso & Fortuin, 2021). Ma et al. (2015) designed a unifying framework for stochastic gradient MCMC whereby all past dynamics-based samplers and their minibatch variants can be cast into it. For the multimodal properties of posterior, Zhang et al. (2020) proposed a cyclical step size schedule in stochastic gradient MCMC for BNNs to improve the sampling efficiency by which they proved the non-asymptotic convergence of their method. On the other hand, Cobb & Jalaian (2021) emphasized the limitations of the use of stochastic gradients in HMC (Shahbaba et al., 2014), and proposed a scaling full HMC for BNNs by a symmetric splitting integration scheme. And Izmailov et al. (2021) also used full-batch HMC to investigate the foundational properties of posterior for BNNs. Alexos et al. (2022) proposed a structured stochastic gradient MCMC that combines structure variational inference and MCMC to improve the slow mixing in high dimensional BNNs.
(a) fSGLD
(b) fSGHMC
(c) FBNN
(d) SGLD
(e) SGHMC
(f) IFBNN
Figure 1: 1-D extrapolation example. The green line is the ground true function, and the blue lines correspond to the mean of samples from posterior predictions. Black dots denote 20 training points; shadow areas represent the predictive standard deviations.
5 Experiments
We evaluate the predictive performance and uncertainty quantification of our functional stochastic gradient MCMC methods on several benchmark tasks, including a synthetic extrapolation example, multivariate regressions on UCI datasets, image classification tasks, and contextual bandits. We compare fSGLD and fSGHMC with naive SGLD, SGHMC, and competing functional variational inference methods to illustrate the sampling efficiency of our functional stochastic gradient MCMC methods.
5.1 Extrapolation on Synthetic Data
To evaluate the fitting ability and uncertainty estimation of our method, we first consider a 1-D oscillation curve extrapolation on a synthetic dataset. For input, we randomly sampled half of 20 observation points from Uniform$(-0.75,-0.25)$, and the other half are from Uniform$(0.25,0.75)$. The output is modeled through the polynomial function: $y=\sin(3\pi x)+0.3\cos(9\pi x)+0.5\sin(7\pi x)+\epsilon$ with noise $\epsilon\sim\mathcal{N}(0,0.5^{2})$. We compared our fSGLD and f SGHMC with naive SGLD, SGHMC, and two representative functional variational methods: FBNN (Sun et al., 2019) and IFBNN (Wu et al., 2023). The model is a fully-connected neural network with two hidden layers. For our functional fSGLD, fSGHMC, and the two functional variational inference methods, we use the same functional GP prior with the RBF kernel pre-trained on the input dataset. There are 40 measurement points for the approximate estimation of the gradient of log functional prior in fSGLD and fSGHMC, which are randomly sampled from training data and an additional 40 inducing points from Uniform $(-1,1)$. For all sampling methods, we use 2000 burn-in iterations and 8000 iterations for 80 samples (to reduce correlations between samples, we draw separated samples every 100 epochs). Functional variational inference methods are trained for 10000 epochs for fair comparison. Results are shown in Figure 1. Figure 1(a) and Figure 1(b) show that fSGLD and fSGHMC can recover the key characteristic of the curve in the observation range and provide reasonable uncertainty estimations in the unseen region. fSGHMC performs slightly better than fSGLD. In contrast, for naive SGMCMC methods, Figure 1(d) illustrates that SGLD underestimates the uncertainty in non-observation areas, especially evident in the leftmost non-observed $[-1,-0.75]$ interval. Meanwhile, Figure 1(e) shows that SGHMC suffers from some unreasonable manifestations of the uncertainty estimation, such as the undeserved uncertainty expansion in the interval $[-0.4,-0.25]$ of the input data range. On the other hand, the rightmost column shows results from two functional variational methods. We found that FBNN in Figure 1(c) is unable to fit the target curve in the training data range $[-0.75,0.25]$ and performs poorly for uncertainty quantification. Figure 1(f) shows that IFBNN also severely underestimates the uncertainty, which is consistent with the findings of Ormerod & Wand (2010) and Zhang et al. (2018) that variational approximation typically underestimates the posterior variance.
5.2 UCI Regression
In this experiment, we evaluate our method for multivariate regression tasks on 7 benchmark UCI datasets: Yacht, Boston, Concrete, Energy, Wine, Kin8nm and Protein. We use a 2- hidden-layer fully connected neural network. For all sampling methods, we run 500 iterations for the burn-in stage and collect 15 samples. Performance is measured via average root mean square error (RMSE) and the test negative log-likelihood (NLL). For each dataset, we construct 10 random 90-to-10 train-test splits and report the mean and standard deviation of RMSE and NLL over the splits in Table 2 and Table 3, respectively. fSGHMC outperforms all baselines for RMSE on all datasets, and fSGLD also achieves better accuracy values than naive SGMCMC methods and functional variational inference, demonstrating superior performance of our functional SGMCMC in predictive accuracy. Moreover, fSGHMC and fSGLD obtain better NLL results on most of the datasets (5/7), showing that our method is also very competitive in uncertainty quantification.
Table 2: The table shows the results of average RMSE for multivariate regression on UCI datasets. We split each dataset randomly into 90% training data and 10% test data, and this process is repeated 10 times to ensure validity.
RMSE
SGLD fSGLD SGHMC fSGHMC FBNN IFBNN
Yacht 1.09 $\pm$ 0.10 0.41 $\pm$ 0.10 1.19 $\pm$ 0.10 0.25 $\pm$ 0.13 1.52 $\pm$ 0.08 1.24 $\pm$ 0.10
Boston 1.23 $\pm$ 0.06 0.36 $\pm$ 0.09 1.31 $\pm$ 0.06 0.24 $\pm$ 0.10 1.68 $\pm$ 0.12 1.44 $\pm$ 0.09
Concrete 1.10 $\pm$ 0.07 0.45 $\pm$ 0.04 1.15 $\pm$ 0.06 0.28 $\pm$ 0.05 1.27 $\pm$ 0.05 1.07 $\pm$ 0.07
Energy 1.04 $\pm$ 0.06 0.24 $\pm$ 0.03 1.07 $\pm$ 0.08 0.18 $\pm$ 0.03 1.35 $\pm$ 0.06 1.19 $\pm$ 0.05
Wine 1.08 $\pm$ 0.09 0.71 $\pm$ 0.05 1.14 $\pm$ 0.09 0.62 $\pm$ 0.04 1.53 $\pm$ 0.05 1.21 $\pm$ 0.05
Kin8nm 1.20 $\pm$ 0.02 0.74 $\pm$ 0.02 1.26 $\pm$ 0.03 0.37 $\pm$ 0.01 1.45 $\pm$ 0.07 1.12 $\pm$ 0.02
Protein 1.12 $\pm$ 0.01 0.84 $\pm$ 0.01 1.21 $\pm$ 0.01 0.83 $\pm$ 0.01 1.50 $\pm$ 0.03 1.16 $\pm$ 0.01
Table 3: The table shows the results of average NLL for multivariate regression on UCI datasets. We split each dataset randomly into 90% training data and 10% test data, and this process is repeated 10 times to ensure validity.
NLL
SGLD fSGLD SGHMC fSGHMC FBNN IFBNN
Yacht -0.37 $\pm$ 0.08 -2.46 $\pm$ 0.28 -2.06 $\pm$ 0.10 -3.37 $\pm$ 0.69 -0.77 $\pm$ 0.86 -1.25 $\pm$ 1.21
Boston -0.60 $\pm$ 0.11 -2.20 $\pm$ 0.20 -2.25 $\pm$ 0.08 -3.26 $\pm$0.33 -1.19 $\pm$ 0.76 0.32 $\pm$ 0.30
Concrete -0.54 $\pm$ 0.05 -1.63 $\pm$ 0.14 -2.51 $\pm$ 0.08 -2.74 $\pm$0.18 -1.00 $\pm$ 0.52 -0.39 $\pm$ 0.33
Energy -0.58 $\pm$ 0.09 -2.25 $\pm$ 0.25 -2.39 $\pm$ 0.09 -4.29 $\pm$0.23 -2.14 $\pm$ 0.47 -1.78 $\pm$ 0.36
Wine -0.90 $\pm$ 0.12 -2.53 $\pm$ 0.15 -2.50 $\pm$ 0.10 -2.34 $\pm$ 0.25 0.52 $\pm$ 0.14 0.26 $\pm$ 0.15
Kin8nm -0.60 $\pm$ 0.04 -2.24 $\pm$ 0.09 -2.29 $\pm$ 0.04 -2.00 $\pm$0.12 -2.45 $\pm$ 0.62 -1.01 $\pm$ 0.14
Protein -0.60 $\pm$ 0.01 -2.10 $\pm$ 0.08 -2.20 $\pm$ 0.01 -2.02 $\pm$0.06 -1.49 $\pm$ 0.24 -2.13 $\pm$ 0.34
Table 4: Image classification and OOD detection performance.
MNIST FMNIST CIFAR10
Model Test Error AUC Test Error AUC Test Error AUC
SGLD 2.73 $\pm$ 0.00 0.859 $\pm$ 0.06 13.92 $\pm$ 0.01 0.633 $\pm$ 0.13 49.56 $\pm$ 0.03 0.669 $\pm$ 0.04
fSGLD 1.89 $\pm$ 0.00 0.960 $\pm$ 0.02 11.66 $\pm$ 0.01 0.849 $\pm$ 0.03 33.57 $\pm$ 0.02 0.672 $\pm$ 0.03
SGHMC 2.97 $\pm$ 0.00 0.853 $\pm$ 0.05 13.72 $\pm$ 0.01 0.589 $\pm$ 0.07 48.13 $\pm$ 0.04 0.639 $\pm$ 0.06
fSGHMC 1.64 $\pm$ 0.00 0.931 $\pm$ 0.02 11.91 $\pm$ 0.00 0.880 $\pm$ 0.03 31.99 $\pm$ 0.01 0.717 $\pm$ 0.03
FBNN 3.99 $\pm$ 0.00 0.801 $\pm$ 0.03 14.36 $\pm$ 0.00 0.814 $\pm$ 0.02 53.71 $\pm$ 0.01 0.612 $\pm$ 0.03
IFBNN 3.64 $\pm$ 0.00 0.949 $\pm$ 0.03 14.15 $\pm$ 0.00 0.838 $\pm$ 0.01 53.38 $\pm$ 0.01 0.616 $\pm$ 0.03
(a) p = 0.4
(b) p = 0.5
(c) p = 0.6
Figure 2: Comparisons of cumulative regrets of fSGLD, SGLD, FBNN, IFBNN, BBB for contextual bandit task on the Mushroom dataset. Lower represents better performance.
5.3 Imagine Classification and OOD Detection
In this section, we investigate the scalability of functional SGMCMC on image classification tasks with high-dimensional inputs. We evaluate the in-distribution predictive performance and out-of-distribution (OOD) detection ability on MNIST (LeCun et al., 2010), FashionMNIST(Xiao et al., 2017) and CIFAR-10(Krizhevsky et al., 2009). We compare our fSGLD and fSGHMC with naive SGLD, SGHMC, functional variational FBNN and IFBNN. We use ResNet-18 architecture (He et al., 2016) for all methods on these three datasets, and the batch size is 125. For all sampling methods, we run for 60 burn-in iterations and collect 10 samples on the MNIST dataset, 100 burn-in iterations and collect 10 samples on FashionMNIST, 80 burn-in iterations and draw 10 samples on CIFAR10. The functional GP prior with the RBF kernel is pre-trained on the test dataset for all functional methods. We randomly sample 40 points from training data in every iteration as the measurement points for the gradient estimation of log functional prior distribution in fSGLD and fSGHMC. We report the test classification errors for predictive performance and the area under the curve (AUC) of OOD detection pairs FashionMNIST/MNIST, MNIST/FashionMNIST and CIFAR10/SVNH based on predictive entropies in Table 4. Our fSGLD and fSGHMC outperform all other methods for both predictive accuracy and OOD detection.
5.4 Contextual Bandits
In many downstream tasks, such as online sequential decision-making scenarios in contextual bandits problems, reliable uncertainty estimation is important to guide the exploration-exploitation process. In such problems, an agent interacts with an unknown environment repeatedly and chooses to take an optimal action in each round of interaction. Thompson sampling (Thompson, 1933; Russo & Van Roy, 2016) is a widely used algorithm for exploration strategies in contextual bandits, it (i) first samples a current posterior to obtain a model configuration; (ii) then adaptively chooses an optimal action based on the current context under the sampled model configuration and observes the corresponding reward; (iii) updates the posterior based on the context, action and reward tuples in this interaction round.
This section evaluates the ability to guide exploration on the UCI Mushroom dataset with 8124 instances. In each instance, the mushroom has 22 features and is labeled as edible or poisonous. An agent can observe mushroom features as the context in each interaction and choose to eat or reject a mushroom. We compare our fSGLD with several baselines, including naive SGLD, functional variational FBNN and IFBNN, parametric-space variational method Bayes By Backprop (BBB) (Blundell et al., 2015). We follow the basic settings by Blundell et al. (2015) and consider three different reward patterns: if the agent eats an edible mushroom, it receives a reward of 5 and a reward of 0 if the agent chooses to reject. On the other hand, if the agent eats a poisonous mushroom, it receives a reward of -35 with probabilities of 0.4, 0.5, and 0.6, respectively, for three different patterns. If the agent chooses to reject a poisonous mushroom, it receives a reward of 0. Suppose an oracle always chooses to eat an edible mushroom and reject the poisonous mushroom. The cumulative regrets with respect to the reward achieved by the oracle measure the exploration-exploitation ability of an agent. We run 10000 iterations for all methods and the minibatch size is 64. Figure 2 shows the cumulative regrets for all methods in three reward patterns. Our fSGLD consistently outperforms other inference methods, indicating that our functional SGMCMC can provide reliable uncertainty estimation in decision-making tasks.
6 Discussion and Conclusions
In this paper, we have lifted the Itô diffusion of network weights in BNNs to function space and proposed a new functional SGMCMC scheme for BNNs. We have designed the functional Langevin dynamics and Hamiltonian dynamics for the function mapping defined by BNNs directly, which can incorporate informative functional prior into posterior inference. Moreover, we proved that the stationary distribution of our functional SGMCMC is the posterior distribution over functions. Empirically, we demonstrated that our functional SGMCMC approaches can leverage functional priors to yield high predictive performance and principled uncertainty quantification on several benchmark tasks.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Further Background
Itô Lemma (Itô, 1951) is a fundamental result in stochastic calculus to find the differential of a function of a stochastic process, which serves as the stochastic calculus counterpart of the chain rule. For an Itô diffusion, let $f(X_{t})$ be an arbitrary twice differentiable scalar function of real variables $X_{t}$, then the differential of $f(X_{t})$ can be derived from the Taylor series expansion of the function as
$\mathrm{d}f(X_{t})=$ $\left(\mu(X_{t})\frac{\partial f}{\partial x}+\frac{\sigma^{2}(X_{t})}{2}\frac{\partial^{2}f}{\partial x^{2}}\right)\mathrm{d}t+\sigma(X_{t})\frac{\partial f}{\partial x}\mathrm{d}B_{t},$ (14)
which implies that $f(X_{t})$ is itself an Itô diffusion (Brzeźniak et al., 2008).
Appendix B Theoretical Proof
The Fokker–Planck equation (Risken, 1996) for the corresponding evolution of the probability density $p(x,t)$ of a diffusion process $X_{t}$ is as
$\frac{\partial p(x,t)}{\partial t}=-\sum_{i}\frac{\partial}{\partial x_{i}}\left[\mu_{i}(X_{t})p(x,t)\right]+\sum_{i,j}\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}\left[D_{i,j}(x,t)p(x,t)\right],$ (15)
where $D_{i,j}(x,t)=\frac{1}{2}\sigma(X_{t})\sigma(X_{t})^{T}$. When the Fokker–Planck equation is equal to zero, it means that the diffusion process has reached a stationary state and $p(x,t)$ is equal to the stationary distribution $\pi(X)$. Then, simulating this diffusion process is equivalent to sampling from the stationary distribution.
The proof of Proposition 3.1 and Proposition 3.2 will be concise based on the general framework for dynamics-based sampling methods proposed by Ma et al. (2015). They proved that if $\mu(X_{t})$ and $\sigma(X_{t})$ are restricted to the following form
$\mu(X)=$ $-[\mathbf{D}(X)+\mathbf{Q}(X)]\nabla G(X)+\Gamma(X),$ (16)
$\Gamma(X)=$ $\sum\frac{\partial}{\partial X}(\mathbf{D}_{ij}(X)+\mathbf{Q}_{ij}(X)),$
$\sigma(X)=$ $\sqrt{2\mathbf{D}(X)},$
where $G(\cdot)$ is the potential energy function, $\mathbf{D}(\cdot)$ is a positive semidefinite matrix, and $\mathbf{Q}(\cdot)$ is a skew-symmetric curl matrix, then the Fokker–Planck equation can be transformed into a more compact form (Yin & Ao, 2006; Shi et al., 2012):
$\frac{\partial p(x,t)}{\partial t}=\nabla^{T}\cdot([\mathbf{D}(X)+\mathbf{Q}(X)][p(x,t)\nabla G(X)+\nabla p(x,t)]).$ (17)
Then, it is easy to verify that the stationary distribution of the diffusion process is exactly as $\pi(X)\propto\exp(-G(X))$.
For the proof of Proposition 3.1, it is clear that the functional Langevin dynamics in Equation 6 can be cast into the above general framework as
$\mathrm{d}f_{t}(\cdot;\mathbf{w})$ $=[-(\nabla_{\mathbf{w}}f)^{T}(\nabla_{\mathbf{w}}f)\nabla_{f}U(f_{t})+H_{\mathbf{w}}f]\mathrm{d}t+\sqrt{2}(\nabla_{\mathbf{w}}f)^{T}\mathrm{d}B_{t}$ (18)
$=\left(-[\mathbf{D}(f)+\mathbf{Q}(f)]\nabla_{f}U(f_{t})+\Gamma(f)\right)\mathrm{d}t+\sqrt{2\mathbf{D}(f)}\mathrm{d}B_{t},$
where $\nabla_{f}U(f)=-\nabla_{f}\log p(\mathcal{D}|f)-\nabla_{f}\log p(f)$, $\mathbf{D}(f)=(\nabla_{\mathbf{w}}f)^{2}$, $\mathbf{Q}(f)=\mathbf{0}$, $\Gamma(f)=H_{\mathbf{w}}f$. Therefore, the Fokker-Planck equation of the probability density $p(f,t)$ for $f_{t}(\cdot;\mathbf{w})$ then can be derived as
$\frac{\partial p(f,t)}{\partial t}=\nabla^{T}\cdot\left([\mathbf{D}(f)+\mathbf{Q}(f)][p(f,t)\nabla_{f}U(f)+\nabla p(f,t)]\right),$ (19)
from which the stationary distribution of $p(f,t)$ can be verified as $\exp(-U(f))=p(f|\mathcal{D})$, that is, the target posterior distribution over functions.
For the proof of Proposition 3.2, the functional Hamiltonian dynamics defined in Equation 10 can be written as
$\mathrm{d}\left[\begin{array}[]{l}f_{t}(\cdot;\mathbf{w})\\ g_{t}(\cdot;\mathbf{z})\end{array}\right]$ $=-\left[\begin{array}[]{cc}0&-(\nabla_{\mathbf{w}}f)^{T}\nabla_{\mathbf{z}}g\\ (\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{w}}f&0\end{array}\right]\left[\begin{array}[]{c}\nabla_{f}U(f_{t})\\ -\nabla_{g}\log p(g_{t})\end{array}\right]\mathrm{d}t$ (20)
$=-\mathbf{Q}(f,g)\nabla H(f_{t},g_{t})\mathrm{d}t,$
where $\mathbf{Q}(f,g)$ is s a skew-symmetric curl matrix, $\mathbf{D}(f,g)=\mathbf{0}$. Therefore, the stationary distributions of the functional Hamiltonian dynamics defined in Equation 10 is $\pi(f,g)\propto\exp(H(f,g))$, and simply (marginalize) discard the auxiliary variables can obtain the target true posterior $p(f|\mathcal{D})$.
For the stochastic gradient version, the functional stochastic gradient Hamiltonian dynamics (fSGHMC) for the discretization update rule for samples of $\mathbf{w}$ and $\mathbf{z}$ defined in Equation 12 is defined as follows:
$\mathrm{d}\left[\begin{array}[]{l}f_{t}(\cdot;\mathbf{w})\\ g_{t}(\cdot;\mathbf{z})\end{array}\right]$ $=-\left[\begin{array}[]{cc}0&-(\nabla_{\mathbf{w}}f)^{T}\nabla_{\mathbf{z}}g\\ (\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{w}}f&C(\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{z}}g\end{array}\right]\left[\begin{array}[]{c}\nabla_{f}\tilde{U}(f_{t})\\ -\nabla_{g}\log p(g_{t})\end{array}\right]\mathrm{d}t+\left[\begin{array}[]{cc}\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\sqrt{2C}(\nabla_{\mathbf{z}}g)^{T}\end{array}\right]\mathrm{d}t,$ (21)
where $\mathbf{D}(f,g)=\left(\begin{array}[]{cc}\mathbf{0}&\mathbf{0}\\ \mathbf{0}&C(\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{z}}g\end{array}\right)$ and $\mathbf{Q}(f,g)=\left(\begin{array}[]{cc}\mathbf{0}&-(\nabla_{\mathbf{w}}f)^{T}\nabla_{\mathbf{z}}g\\ (\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{w}}f&\mathbf{0}\end{array}\right)$ is the same as that in full-batch functional Hamiltonian dynamics. Then, the (marginal) stationary distribution of such functional stochastic gradient Hamiltonian is still the target functional posterior $p(f|\mathcal{D})$.
Moreover, the comparisons of matrix $\mathbf{D}(\cdot)$ and $\mathbf{Q}(\cdot)$ for naive SGMCMC methods and our functional SGMCMC are shown in Table 5.
Table 5: General framework of naive SGMCMC and functional SGMCMC algorithms.
Dynamics $X$ $G(X)$ $\mathbf{Q}(X)$ $\mathbf{D}(X)$
SGLD $\mathbf{w}$ $\tilde{U}(\mathbf{w})$ $\mathbf{0}$ $\mathbf{I}_{d}$
fSGLD $f$ $\tilde{U}(f)$ $\mathbf{0}$ $(\nabla_{\mathbf{w}}f)^{T}(\nabla_{\mathbf{w}}f)$
SGHMC $(\mathbf{w},\mathbf{z})$ $\tilde{H}(\mathbf{w},\mathbf{z})$ $\left(\begin{array}[]{cc}\mathbf{0}&-\mathbf{I}_{d}\\ \mathbf{I}_{d}&\mathbf{0}\end{array}\right)$ $\left(\begin{array}[]{cc}\mathbf{0}&\mathbf{0}\\ \mathbf{0}&C\end{array}\right)$
fSGHMC $(f,g)$ $\tilde{H}(f,g)$   $\left(\begin{array}[]{cc}\mathbf{0}&-(\nabla_{\mathbf{w}}f)^{T}\nabla_{\mathbf{z}}g\\ (\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{w}}f&\mathbf{0}\end{array}\right)$ $\left(\begin{array}[]{cc}\mathbf{0}&\mathbf{0}\\ \mathbf{0}&C(\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{z}}g\end{array}\right)$
Appendix C Pseudocode for Functional SGMCMC.
Algorithm 1 presents the pseudocode for fSGLD. And the pseudocode for fSGHMC is shown in Algorithm 2.
Algorithm 1 Functional SGLD (fSGLD)
  Input: Dataset $\mathcal{D}$, pre-trained GP prior $p(f)$, initialized $\mathbf{w}_{0}\sim\mathcal{N}(0,I)$, step size $\epsilon_{t}$, noise $\eta_{t}\sim\mathcal{N}(0,1)$, number of burn-in iterations $N_{b}$, number of sample $K$
  ##\## Burn-in stage :
  for $t=0$ to $N_{b}$ do
     draw measurement set $\mathbf{X}_{\mathcal{M}}$;
     update $\mathbf{w}_{t+1}\leftarrow\mathbf{w}_{t}-\epsilon_{t}\nabla_{\mathbf{w}}\tilde{U}(f_{t}))+\sqrt{2\epsilon_{t}}\eta_{t}$ using Equation 9;
  end for
  ##\## Sampling stage :
  $S\leftarrow\varnothing$;
  $\mathbf{w}_{0}\leftarrow\mathbf{w}_{N_{b}}$;
  for $t=0$ to $K$ do
     draw measurement set $\mathbf{X}_{\mathcal{M}}$;
     $\mathbf{w}_{t+1}\leftarrow\mathbf{w}_{t}-\epsilon_{t}\nabla_{\mathbf{w}}\tilde{U}(f_{t}))+\sqrt{2\epsilon_{t}}\eta_{t}$ using Equation 9;
     $S\leftarrow S\cup\left\{\mathbf{w}_{t+1}\right\}$;
  end for
Algorithm 2 Functional SGHMC (fSGHMC)
  Input: Dataset $\mathcal{D}$, pre-trained GP prior $p(f)$, initialized $\mathbf{w}_{0}\sim\mathcal{N}(0,I)$, $\mathbf{z}_{0}\sim\mathcal{N}(0,M)$, step size $\epsilon_{t}$, noise $\eta_{t}\sim\mathcal{N}(0,1)$, number of burn-in iterations $N_{b}$, leapfrog steps $m$, number of sample $K$
  ##\## Burn-in stage :
  for $t=0$ to $N_{b}$ do
     resample momentum $\mathbf{z}_{t}\sim\mathcal{N}(0,M)$
     $(\mathbf{w}_{0},\mathbf{z}_{0})\leftarrow(\mathbf{w}_{t},\mathbf{z}_{t})$;
     for $i=1$ to $m$ do
        draw measurement set $\mathbf{X}_{\mathcal{M}}$;
        update $(\mathbf{w}_{i},\mathbf{z}_{i})$ using Equation 12;
     end for
     $(\mathbf{w}_{t+1},\mathbf{z}_{t+1})\leftarrow(\mathbf{w}_{m},\mathbf{z}_{m})$
  end for
  ##\## Sampling stage :
  $S\leftarrow\varnothing$;
  $\mathbf{w}_{0}\leftarrow\mathbf{w}_{N_{b}}$;
  for $t=0$ to $K$ do
     resample momentum $\mathbf{z}_{t}\sim\mathcal{N}(0,M)$
     $(\mathbf{w}_{0},\mathbf{z}_{0})\leftarrow(\mathbf{w}_{t},\mathbf{z}_{t})$;
     for $i=1$ to $m$ do
        draw measurement set $\mathbf{X}_{\mathcal{M}}$;
        update $(\mathbf{w}_{i},\mathbf{z}_{i})$ using Equation 12;
     end for
     $(\mathbf{w}_{t+1},\mathbf{z}_{t+1})\leftarrow(\mathbf{w}_{m},\mathbf{z}_{m})$
     $S\leftarrow S\cup\left\{\mathbf{w}_{t+1}\right\}$;
  end for
Appendix D Ablation Study
D.1 The effects of the sample size for naive SGMCMC and functional SGMCMC
In this section, we investigate the effect of sample size on the results of naive SGMCMC and our functional SGMCMC by varying the number of samples. For the 1-D extrapolation example in Section 5.1. We consider three different sample sizes for SGLD, SGHMC, fSGLD, and fSGHMC: 10, 80, and 200, respectively. The extrapolation results are shown in Figure 3. We can see that no matter the sample size, there is almost no difference in the fitting effect and uncertainty estimation for both naive SGMCMC and our functional SGMCMC, which indicates that the sample size has almost no effect on our experimental results.
(a) fSGLD-10
(b) fSGLD-80
(c) fSGLD-200
(d) SGLD-10
(e) SGLD-80
(f) SGLD-200
(g) fSGHMC-10
(h) fSGHMC-80
(i) fSGHMC-200
(j) SGHMC-10
(k) SGHMC-80
(l) SGHMC-200
Figure 3: The effect of sample size on 1-D extrapolation example for fSGLD, SGLD, fSGHMC, SGHMC. The number after the short line in each subheading represents the sample size.
For the UCI regressions in Section 5.2, the sample size is 15 for all naive SGMCMC and functional SGMCMC methods. We now consider a bigger sample size of 150 for all methods on the Yacht and Boston datasets. The RMSE and NLL results for the naive SGMCMC and functional SGMCMC are shown in Table 6 and Table 7, respectively. We randomly split each dataset into 90% training data and 10% test data, which is repeated 5 times to ensure validity. These two results have only very slight fluctuations compared to the results in Table 2 and Table 3, which still illustrates that our results are almost independent of the sample size. Our functional fSGLD and fSGHMC consistently demonstrate superior performance.
Table 6: The table shows the average RMSE results for sample size effects (150 samples).
RMSE
SGLD fSGLD SGHMC fSGHMC
Yacht 1.095 $\pm$ 0.128 0.409 $\pm$ 0.112 1.187 $\pm$ 0.141 0.246 $\pm$ 0.142
Boston 1.248 $\pm$ 0.072 0.392 $\pm$ 0.112 1.332 $\pm$ 0.068 0.270 $\pm$ 0.135
Table 7: The table shows the average NLL results for sample size effects (150 samples).
NLL
SGLD fSGLD SGHMC fSGHMC
Yacht -0.352 $\pm$ 0.073 -2.332 $\pm$ 0.493 -2.300 $\pm$ 0.181 -3.406 $\pm$ 0.517
Boston -0.509 $\pm$ 0.129 -2.068 $\pm$ 0.159 -2.352 $\pm$ 0.100 -3.060 $\pm$ 0.284
D.2 The effects of the number of measurement points on the gradient estimation of functional prior
In the 1-D extrapolation example in Section 5.1, we sampled 40 measurement points from 20 training data and an additional 40 inducing points from Uniform $(-1,1)$ to estimate the gradient of log functional prior for fSGLD and fSGHMC. We now investigate the effects of the number of measurement points on the approximate estimation of the gradient of the log functional prior in functional SGMCMC. We consider two other cases for the number of measurement points: 10 measurement points sampled from 20 training data and an additional 40 inducing points from Uniform $(-1,1)$; and 400 measurement points sampled from 20 training data and an additional 1000 inducing points from Uniform $(-1,1)$. The comparison results are shown in Figure 4. We can see that there is almost no difference in the prediction performance for the three different numbers of measurement points. The uncertainty estimations also fluctuate only slightly. Overall, we can conclude that the number of measurement points for gradient estimation of functional prior does not significantly impact the results of our methods on 1-D extrapolation example.
(a) fSGLD-m10
(b) fSGLD-m40
(c) fSGLD-m400
(d) fSGHMC-m10
(e) fSGHMC-m40
(f) fSGHMC-m400
Figure 4: The effects of the number of measurement points on the gradient estimation of functional prior for fSGLD and fSGHMC. The number after the short line in each subheading represents the number of measurement points.
In the UCI regressions in Section 5.2, for datasets with sample sizes less than 1000, we estimate the gradient of the functional prior for fSGLD and fSGHMC using the training data as the finite measurement points. For datasets with sample sizes larger than 1000, we randomly sample 1000 data from the training data as measurement points. We now consider using only 40 measurement points on the Yacht and Boston datasets. The results of RMSE and NLL are shown in Table 8 and Table 9, respectively. Compared to the results in Table 2 and Table 3, the RMSE results are slightly worse for both fSGLD and fSGHMC. The NLL results for fSGLD get slightly worse on the Yacht dataset and remain essentially unchanged on the Boston dataset. The NLL results for fSGHMC are slightly worse on both datasets but still significantly outperforms all other methods. Overall, for the UCI regressions, the more measurement points we used, the better the performance of our methods, probably because more measurement points are more accurate for the gradient estimation of the functional prior.
Table 8: The table shows the average RMSE results for the effects of the number of measurement points.
RMSE
fSGLD fSGHMC
Yacht 0.496 $\pm$ 0.072 0.297 $\pm$ 0.111
Boston 0.448 $\pm$ 0.098 0.290 $\pm$ 0.126
Table 9: The table shows the average NLL results for the effects of the number of measurement points.
NLL
fSGLD fSGHMC
Yacht -2.327 $\pm$ 0.564 -2.738 $\pm$ 0.428
Boston -2.277 $\pm$ 0.089 -2.607 $\pm$ 0.329
D.3 The effects of the kernel function on the functional GP prior
(a) fSGLD-RBF
(b) fSGLD-Linear
(c) fSGLD-Matern
(d) fSGHMC-RBF
(e) fSGHMC-Linear
(f) fSGHMC-Matern
Figure 5: The effects of the kernel function on the functional GP prior. The text after the short line in each subheading represents the corresponding kernel functions.
Functional Gaussian processes (GPs) priors are able to encode prior knowledge about function properties (e.g., periodicity and smoothness) through corresponding kernel functions. We used RBF kernel in the 1-D extrapolation experiment, which is suited for modeling polynomial functions. We now consider two other kernel functions of GP prior to our functional SGMCMC: Matern kernel and Linear kernel (unsuitable for modeling polynomial oscillatory curves). The results are shown in Figure 5. The results from the Linear kernel in Figure 5(b) and Figure 5(e) for fSGLD and fSGHMC respectively show that the fitting performance deteriorates significantly in the observation region and also demonstrate a strong linear trend (converging to a horizontal line) in the middle unseen region $[-0.25,0.25]$. The uncertainty intervals in the leftmost $[-1,-0.75]$ and rightmost $[0.75,1]$ non-observation regions deviate far from the true function. On the contrary, the results of the Matern kernel are hardly different from those of the RBF kernel, due to the fact that it also has a strong ability to describe polynomial functions.